OpenStudy (anonymous):

Add and subtract as directed. In your answer place a" ^" before the exponent. Example 2x^4 is 2x^4.) a-7/3+ 5a-3/6-2a+5/4=_______

OpenStudy (anonymous):

is it a^-7 or a-7

OpenStudy (anonymous):

it`s a-7

OpenStudy (radar):

is it \[a-7 \over 3\]or\[a-7\over 3+5a\]

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

16a-38/24 is the answer

OpenStudy (anonymous):

first you'll make the denominator same

OpenStudy (anonymous):

to same them take LCM of all denominators

OpenStudy (anonymous):

you have to make all denominators 24

OpenStudy (anonymous):

so multiply the first term by 8 second term by 4 and third term by 6

OpenStudy (anonymous):

after denominator is same then its simple addition and subtraction

OpenStudy (radar):

O.K.\[(a-7)/(3) + 5a - 3/6 -2a + 5/4=\] First combine simlar terms getting (a-7)/3 +3a+3/4=

OpenStudy (radar):

the lowest common denominator LCD is 12 and the fractions now become (4a-28)/12 +60a/12 +9/12= (64a-19)/12 Assuming I have correctly copied it.

OpenStudy (radar):

I think I may have erred when converting to LCD try again (4a-28)/12 +36a/12 + 9/12= 40a/12 - 19/12 = 10a/3 -19/12=

OpenStudy (anonymous):

Thank you @annas :) and thank you radar!

OpenStudy (anonymous):

no problem maria !